• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name jpchen

    Log entry time 16:19:44 on March 7,2012

    Entry number 364295

    keyword=Day shift summary

    Shift Summary

    Shift workers: J. P. Chen (SL), K. Myers, Y. Zhao.

    8:00 There was problem over night that not be able to get beam through the chicane. After 8:00 am discussion, Yves had a long list of things to check. Hall is in controlled access.
    9:00 Gte Hall A back to beam permit for harp scan and some beam tuning/check.
    10:30-11:30 beam tuning and check. Could not do harp.
    12:00-16:00 controlled access to check magnets (FZ1/FZ2, BPM position, harp etc.
    It was found that the likely reason for the problem was that the FZ2 off by 5% (Set/readback 401A, measured at the leads 423 A.